OpenStudy (anonymous):

the perimeter of a college athletic field is 96 meters and the length is 12m more than the width. find the width and length?

OpenStudy (anonymous):

P = 2(L +B) 96 = 2(x +[x+12]) 96 = 2(2x+12) 96=4X +24 -24 -24 72 = 4X divide by 4 18 = x (which is the width) 30 = x+12 (which is the length)

OpenStudy (anonymous):

i figure it was 18 for W. thx for the length.